OpenStudy (anonymous):

Which of the following is the simplest form of this expression? 5sqr(a^4)/3sqr(a^2) by sqr i mean square root. A) a 2/15 B) 1/a 7/4 c) a 7/4 d) 1/ a 2/15

Nnesha (nnesha):

\[\huge\rm \frac{ 4\sqrt{a^4} }{ 3 \sqrt{a^2}}\] like this ?

OpenStudy (anonymous):

yes

OpenStudy (triciaal):

** 5 not 4

OpenStudy (anonymous):

yes

Nnesha (nnesha):

\[\huge\rm \frac{ 5\sqrt{a^4} }{ 3 \sqrt{a^2}}\] typo thanks triciaal

Nnesha (nnesha):

so you can convert square root to 1/2 exponent \[\huge\rm \sqrt[n]{x^m} = x^\frac{ m }{ n }\]

OpenStudy (anonymous):

and

Nnesha (nnesha):

do it and let me know what you get \[\huge\rm \sqrt{a^4} = a^\frac{ ? }{ ?? }\]

OpenStudy (anonymous):

a 5/4

Nnesha (nnesha):

square root so n=2

OpenStudy (triciaal):

above is correct you can also use (rt a ) as a single entity when you divide numbers or variable with the same base subtract the exponents

Nnesha (nnesha):